OpenStudy (anonymous):

Graph the following function using transformations. y=1/4(x+)^2 show all work

geerky42 (geerky42):

"(x+)^2" ???

OpenStudy (anonymous):

y=1/4(x+4)^2

geerky42 (geerky42):

Like this? \( \Large y = \dfrac{1}{4}(x+4)^2 \)

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

can you help me?

OpenStudy (cwrw238):

start with the graph of y = x^2 the + 4 moves the whole graph 4 units to the left as for the 1/4 y = af(x) is a stretch factor a parallel to the y-axis

OpenStudy (anonymous):

what do you mean by y=af(x)

OpenStudy (cwrw238):

a multiplies f(x) in this case its 1/4 so here we have a stretch factor 1/4 which is really a 'squash' every value of y in y=x^2 is multiplied by a quarter

OpenStudy (cwrw238):

the graph will look like this http://www.wolframalpha.com/input/?i=y+%3D+0.25%28x+%2B+4%29%5E2

OpenStudy (anonymous):

thank you

OpenStudy (cwrw238):

yw